# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ II System Model and Problem Formulation
II-A Channel Model
In this paper, we assume that the movable regions at both the users and the BS are much smaller than the signal propagation distance, ensuring the far-field condition. Therefore, the angle of departure (AoD), the angle of arrival (AoA), and the amplitude of the complex gain for each channel path remain unchanged regardless of the FA positions. In contrast, the phase of each channel path depends on the FA positions and varies accordingly [24].
@@ II-A Channel Model
Let $L_{k}^{t}$ and $L_{k}^{r}$, $1\leq k\leq K$, denote the numbers of transmit and receive channel paths between the BS and user $k$, respectively. The elevation and azimuth AoDs of the $i$-th transmit path are denoted by $\theta_{ki}^{t}$ and $\phi_{ki}^{t}$, where $1\leq i\leq L_{k}^{t}$. Similarly, the elevation and azimuth AoAs of the $j$-th receive path are denoted by $\theta_{kj}^{r}$ and $\phi_{kj}^{r}$, where $1\leq j\leq L_{k}^{r}$. Therefore, the steering vectors corresponding to the directions of the $i$-th transmit and $j$-th receive paths are given by
$\mathbf{g}_{ki}^{t}$ $=\left[\cos\theta_{ki}^{t}\cos\phi_{ki}^{t},\cos\theta_{ki}^{t}\sin\phi_{ki}^{t},\sin\theta_{ki}^{t}\right]^{\mathsf{T}},$ (2)
$\mathbf{f}_{kj}^{r}$ $=\left[\cos\theta_{kj}^{r}\cos\phi_{kj}^{r},\cos\theta_{kj}^{r}\sin\phi_{kj}^{r},\sin\theta_{kj}^{r}\right]^{\mathsf{T}},$ (3)
with $1\leq k\leq K$, $1\leq i\leq L_{k}^{t}$, and $1\leq j\leq L_{k}^{r}$, respectively. Then, for the $i$-th transmit channel path from the BS to user $k$, the distance difference between the path originating from the $m$-th BS antenna position $\mathbf{t}_{m}$ and that from the origin of the BS coordinate system $\mathbf{o}$ is given by
$\rho_{ki}^{t}(\mathbf{t}_{m})\triangleq\mathbf{g}_{ki}^{t}\left(\mathbf{t}_{m}-\mathbf{o}\right)$ (4)
$=$ $x_{m}^{t}\cos\theta_{ki}^{t}\cos\phi_{ki}^{t}+y_{m}^{t}\cos\theta_{ki}^{t}\sin\phi_{ki}^{t}+z_{m}^{t}\sin\theta_{ki}^{t}.$
@@
Similarly, for the $j$-th receive channel path from the BS to user $k$, the distance difference between the path originating from the $n$-th receive antenna position $\mathbf{r}_{kn}$ and that from the origin of user $k$’s coordinate system $\mathbf{o}_{k}$ is given by
$\rho_{kj}^{r}(\mathbf{r}_{kn})\triangleq\mathbf{f}_{kj}^{r}\left(\mathbf{r}_{kn}-\mathbf{o}_{k}\right)$ (5)
$=$ $x_{kn}^{r}\cos\theta_{kj}^{r}\cos\phi_{kj}^{r}+y_{kn}^{r}\cos\theta_{kj}^{r}\sin\phi_{kj}^{r}+z_{kn}^{r}\sin\theta_{kj}^{r}.$
@@
Let $\lambda$ denote the carrier wavelength. For the $i$-th transmit (and $j$-th receive) channel paths from the BS to user $k$, the phase differences between the path originating from $\mathbf{t}_{m}$ and that from $\mathbf{o}_{m}$ (and between the path originating from $\mathbf{r}_{kn}$ and that from $\mathbf{o}_{kn}$) are given by $2\pi\rho_{ki}^{t}(\mathbf{t}_{m})/\lambda$ and $2\pi\rho_{kj}^{r}(\mathbf{r}_{kn})/\lambda$, respectively. Hence, the transmit and receive field-response vectors (FRVs) for the channel between the BS and user $k$ are given by [23, 12]
@@
$\mathbf{g}_{k}(\mathbf{t}_{m})$ $\triangleq\left[\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,1}^{t}(\mathbf{t}_{m})},\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,2}^{t}(\mathbf{t}_{m})},\cdots,\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,L_{k}^{t}}^{t}(\mathbf{t}_{m})}\right]^{\mathsf{T}},$ (6)
@@
$\mathbf{f}_{k}(\mathbf{r}_{kn})$ $\triangleq\left[\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,1}^{r}(\mathbf{r}_{kn})},\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,2}^{t}(\mathbf{r}_{kn})},\cdots,\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,L_{k}^{r}}^{t}(\mathbf{r}_{kn})}\right]^{\mathsf{T}},$ (7)
with $1\leq k\leq K$, $1\leq m\leq M$, and $1\leq n\leq N$.
@@
Besides, we define the path-response matrix (PRM) $\mathbf{\Sigma}_{k}\in\mathbb{C}^{L_{k}^{r}\times L_{k}^{t}}$ as the response between each pair of transmit and receive channel paths from the BS to user $k$. The entry $[\mathbf{\Sigma}_{k}]_{ji}$ represents the response between the $i$-th transmit and $j$-th receive paths for user $k$. Thus, $\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)$ is given by
$\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)=\mathbf{F}_{k}^{\mathsf{H}}\left(\mathbf{\mathbf{R}_{k}}\right)\mathbf{\Sigma}_{k}\mathbf{G}_{k}\left(\mathbf{T}\right),\vspace{0.2cm}$ (8)
@@
where $\mathbf{F}_{k}\left(\mathbf{R}_{k}\right)=\left[\mathbf{f}_{k}(\mathbf{r}_{k,1}),\cdots,\mathbf{f}_{k}(\mathbf{r}_{kn})\right]\in\mathbb{C}^{L_{k}^{r}\times N}$ and $\mathbf{G}_{k}\left(\mathbf{T}\right)=\left[\mathbf{g}_{k}(\mathbf{t}_{1}),\cdots,\mathbf{g}_{k}(\mathbf{t}_{M})\right]\in\mathbb{C}^{L_{k}^{t}\times M}$ denote the field response matrices (FRMs) of all the receive FAs at user $k$ and those at the BS, respectively.
II-B Problem Formulation
@@
$f_{\rm Lag}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}}\right)=$ $\sum_{k=1}^{K}\alpha_{k}\left(\log\det\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)-\mathtt{tr}\left(\mathbf{\Gamma}_{k}\right)+\mathtt{tr}\left(\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)\mathbf{W}_{k}^{\mathsf{H}}\mathbf{H}_{k}^{\mathsf{H}}\left(\mathbf{T},\mathbf{R}_{k}\right)\right.\right.$ (14)
×$\left.\left.\times\left(\sum\nolimits_{j=1}^{K}\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{j}\right)\mathbf{W}_{j}\mathbf{W}_{j}^{\mathsf{H}}\mathbf{H}_{k}^{\mathsf{H}}\left(\mathbf{T},\mathbf{R}_{j}\right)+\sigma_{k}^{2}\mathbf{I}\right)^{-1}\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)\mathbf{W}_{k}\right)\right).$
$f_{\rm Quad}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}\right)=$ $\sum_{k=1}^{K}\left(\alpha_{k}\log\det\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)-\alpha_{k}\mathtt{tr}\left(\mathbf{\Gamma}_{k}\right)+\mathtt{tr}\left(\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)\left(\sqrt{\alpha_{k}}\mathbf{W}_{k}^{\mathsf{H}}\mathbf{H}_{k}^{\mathsf{H}}\left(\mathbf{T},\mathbf{R}_{k}\right)\mathbf{\Phi}_{k}\right.\right.\right.$ (16)
@@
$\underset{\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}}}{\max}$ $R$ (12a)
s. t. $\sum_{k=1}^{K}\mathtt{tr}\left(\mathbf{W}_{k}\mathbf{W}_{k}^{\mathsf{H}}\right)\leq P_{\max},$ (12b)
$\mathbf{t}_{m}\in\mathcal{C}_{m}^{t},\ \forall m,$ (12c)
$\mathbf{r}_{kn}\in\mathcal{C}_{kn}^{r},\ \forall{kn},$ (12d)
[…]
Here, the constraint (12b) denotes the total transmit power constraint, where $P_{\max}$ is the total transmit power budget at the BS. Constraints (12c) and (12d) guarantee the FAs at the BS and users remain within the movable regions. Constraints (12e) and (12f) prevent coupling effects between any pair of FAs at the BS and users, respectively. The problem (12) is difficult to solve because the objective function (12a) is highly nonlinear and non-concave with respect to (w.r.t.) the beamforming matrices $\underline{\mathbf{W}}$ and the FA positions $\mathbf{T}$ and $\underline{\mathbf{R}}$. Additionally, the optimization variables are highly coupled, making the problem more intractable.
III Block Coordinate Ascent (BCA)-Based Algorithm
In this section, we propose a BCA-based algorithm to the problem (12). First, we employ the FP method to decouple the variables in the problem (12). Then, the MM algorithm is utilized to address the non-concave optimization of FA positions.
III-A Problem Reformulation
To solve the complicated problem (12), we first reformulate it using the widely adopted FP method [27]. Specifically, since the objective function (12a) is a sum-of-functions-of-matrix-ratios, we apply the matrix FP framework developed in [34], detailed as follows.
First, applying the matrix Lagrangian dual transform [34, Theorem 2] to the problem (12) allows us to extract the ratios from the logarithms in (10). Therefore, the problem (12) can be reformulated as
$\underset{\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}}}{\max}$ $f_{\rm Lag}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}}\right)$ (13a)
@@ III-A Problem Reformulation
By applying the matrix quadratic transform [34, Theorem 1] to the problem (13), we can further decouple the ratios in the reformulated objective function (14) and reformulate the problem (13) as
$\underset{\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}}{\max}$ $f_{\rm Quad}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}\right)$ (15a)
@@
III-B The Update Step of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
In this step, we aim to optimize the auxiliary variables $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ with the fixed $\underline{\mathbf{W}}$, $\mathbf{T}$, and $\underline{\mathbf{R}}$. The optimal solution of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ are derived by setting the first-order derivatives of (16) to zero w.r.t. $\mathbf{\Gamma}_{k}$ and $\mathbf{\Phi}_{k}$, respectively. Let $\overline{{\mathbf{W}}}_{k}$, $\overline{\mathbf{T}}$, $\overline{\mathbf{R}}_{k}$, $\overline{\mathbf{\Gamma}}_{k}$, and let $\overline{\mathbf{\Phi}}_{k}$ denote the temporal optimization results obtained by the previous iteration, and let $\overline{\mathbf{M}}_{k}$ denote the temporal interference-plus-noise matrix. The closed-form expressions for the optimal $\mathbf{\Phi}_{k}$ and $\mathbf{\Gamma}_{k}$ are given by [34]
@@ III-B The Update Step of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
III-C The Update Step of $\underline{\mathbf{W}}$
@@
$\mathbf{W}_{k}=$ $\left(\sum\nolimits_{j=1}^{K}\overline{\mathbf{H}}_{j}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\overline{\mathbf{H}}_{j}+\mu\mathbf{I}\right)^{-1}$
×$\times\sqrt{\alpha_{k}}\overline{\mathbf{H}_{k}}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right),$ (20)
where $\mu\geq 0$ is computed via bisection search to ensure that $\overline{\mathbf{W}}$ can satisfy the complementary slackness condition of the power budget constraint (12b[35].
III-D The Update Step of $\mathbf{T}$
$\delta^{\mathtt{Tx}}=\underset{1\leq m\leq M}{\max}\frac{24\pi^{2}}{\lambda^{2}}$ $\sum_{k=1}^{K}L_{k}^{t}\left(\left(\sum_{j=1}^{M}\Big{\lvert}[\hat{\mathbf{W}}]_{mj}\Big{\rvert}+\sqrt{M}\Big{\lVert}[\hat{\mathbf{W}}]_{m}\Big{\rVert}_{2}\right)\lVert\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\rVert_{2}+\sqrt{\frac{\alpha_{k}}{L_{k}^{t}}}\Big{\lVert}[\overline{\mathbf{W}}_{k}]_{m}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right).$ (31)

### v2

@@ III-A Signal Model
III-B Channel Model
@@
Let $L_{k}^{\mathtt{Tx}}$ and $L_{k}^{\mathtt{Rx}}$ denote the numbers of transmit and receive channel paths between the BS and user $k$, respectively. The direction vectors corresponding to the $q$-th transmit and receive paths are given by
$\mathbf{g}_{kq}^{\mathtt{Tx}}$ $=\left[\cos\theta_{kq}^{\mathtt{Tx}}\cos\phi_{kq}^{\mathtt{Tx}},\cos\theta_{kq}^{\mathtt{Tx}}\sin\phi_{kq}^{\mathtt{Tx}},\sin\theta_{kq}^{\mathtt{Tx}}\right]^{\mathsf{T}},$ (2)
$\mathbf{f}_{kq}^{\mathtt{Rx}}$ $=\left[\cos\theta_{kq}^{\mathtt{Rx}}\cos\phi_{kq}^{\mathtt{Rx}},\cos\theta_{kq}^{\mathtt{Rx}}\sin\phi_{kq}^{\mathtt{Rx}},\sin\theta_{kq}^{\mathtt{Rx}}\right]^{\mathsf{T}},$ (3)
@@
where $\theta_{kq}^{\mathtt{Tx}}$ and $\phi_{kq}^{\mathtt{Tx}}$ (and $\theta_{kq}^{\mathtt{Rx}}$ and $\phi_{kq}^{\mathtt{Rx}}$) are the elevation and azimuth AoDs (and AoAs) of the $q$-th path between the BS and user $k$. For the $q$-th transmit (and receive) channel path from the BS to user $k$, the distance difference between the path originating from the $m$-th BS antenna position $\mathbf{t}_{m}$ (and $n$-th user antenna position $\mathbf{r}_{kn}$) and that from the origin of the BS (and user $k$) coordinate system are given by
$\rho_{kq}^{\mathtt{Tx}}(\mathbf{t}_{m})\triangleq\left(\mathbf{g}_{kq}^{\mathtt{Tx}}\right)^{\mathsf{T}}\mathbf{t}_{m},\qquad\rho_{kq}^{\mathtt{Rx}}(\mathbf{r}_{kn})\triangleq\left(\mathbf{f}_{kq}^{\mathtt{Rx}}\right)^{\mathsf{T}}\mathbf{r}_{kn},$ (4)
@@
respectively. The transmit and receive field-response vectors (FRVs) between the BS and user $k$ are given by [23]
@@
$\mathbf{g}_{k}(\mathbf{t}_{m})$ $\triangleq\big[\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k1}^{\mathtt{Tx}}(\mathbf{t}_{m})},\dots,\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,L_{k}^{\mathtt{Tx}}}^{\mathtt{Tx}}(\mathbf{t}_{m})}\big]^{\mathsf{T}},$ (5)
@@
$\mathbf{f}_{k}(\mathbf{r}_{kn})$ $\triangleq\big[\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k1}^{\mathtt{Rx}}(\mathbf{r}_{kn})},\cdots,\mathrm{e}^{\jmath\frac{2\pi}{\lambda}\rho_{k,L_{k}^{\mathtt{Rx}}}^{\mathtt{Rx}}(\mathbf{r}_{kn})}\big]^{\mathsf{T}},$ (6)
@@
respectively, where $\lambda$ denotes the carrier wavelength. By defining the path-response matrix (PRM) $\mathbf{\Sigma}_{k}\in\mathbb{C}^{L_{k}^{\mathtt{Rx}}\times L_{k}^{\mathtt{Tx}}}$ as the response between each pair of transmit and receive channel paths from the BS to user $k$, the channel matrix $\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)$ is given by
$\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)=\mathbf{F}_{k}^{\mathsf{H}}\left(\mathbf{\mathbf{R}_{k}}\right)\mathbf{\Sigma}_{k}\mathbf{G}_{k}\left(\mathbf{T}\right),$ (7)
@@
where $\mathbf{F}_{k}\left(\mathbf{R}_{k}\right)=\left[\mathbf{f}_{k}(\mathbf{r}_{k1}),\dots,\mathbf{f}_{k}(\mathbf{r}_{kn})\right]$ and $\mathbf{G}_{k}\left(\mathbf{T}\right)=\left[\mathbf{g}_{k}(\mathbf{t}_{1}),\dots,\mathbf{g}_{k}(\mathbf{t}_{M})\right]$ denote the field response matrices (FRMs) of all the receive FAs at user $k$ and those at the BS, respectively.
III-C Problem Formulation
@@
[…]
$\left.\left.\times\left(\sum\nolimits_{j=1}^{K}\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)\mathbf{W}_{j}\mathbf{W}_{j}^{\mathsf{H}}\mathbf{H}_{k}^{\mathsf{H}}\left(\mathbf{T},\mathbf{R}_{k}\right)+\sigma_{k}^{2}\mathbf{I}\right)^{-1}\mathbf{H}_{k}\left(\mathbf{T},\mathbf{R}_{k}\right)\mathbf{W}_{k}\right)\right).$
[…]
@@
$\underset{\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}}}{\max}\ \$$R$ (11a)
s. t. $\sum\nolimits_{k=1}^{K}\mathtt{tr}\left(\mathbf{W}_{k}\mathbf{W}_{k}^{\mathsf{H}}\right)\leq P_{\max},$ (11b)
$\mathbf{t}_{m}\in\mathcal{C}_{m}^{\mathtt{Tx}},\ \forall m,$ (11c)
$\mathbf{r}_{kn}\in\mathcal{C}_{kn}^{\mathtt{Rx}},\ \forall{kn},$ (11d)
$\|\mathbf{t}_{m}-\mathbf{t}_{m^{\prime}}\|_{2}\geq D,\ 1\leq m,m^{\prime}\leq M,m\neq m^{\prime},$ (11e)
$\|\mathbf{r}_{kn}-\mathbf{r}_{kn^{\prime}}\|_{2}\geq D,\ \forall k,\ 1\leq n,n^{\prime}\leq N,n\neq n^{\prime}.$ (11f)
Here, the constraint (11b) denotes the total transmit power constraint, where $P_{\max}$ is the total transmit power budget at the BS. Constraints (11c) and (11d) guarantee the FAs at the BS and users remain within the movable regions. Constraints (11e) and (11f) prevent mechanical collision between any pair of FAs at the BS and users, respectively. The problem (11) is difficult to solve because the objective function (11a) is highly non-linear and non-concave w.r.t. the beamforming matrices $\underline{\mathbf{W}}$ and the FA positions $\mathbf{T}$ and $\underline{\mathbf{R}}$. Additionally, the optimization variables are highly coupled, making the problem more intractable.
IV Block Coordinate Ascent (BCA)-Based Algorithm
In this section, we propose a BCA-based algorithm for the problem (11). First, we employ the matrix FP method to decouple the variables in the problem (11). Then, the MM algorithm is utilized to address the non-convex optimization of FA positions.
IV-A Problem Reformulation
To solve the complicated problem (11), we first reformulate it using the matrix FP method [27]. Specifically, since the objective function (11a) is a sum-of-functions-of-matrix-ratios, we apply the matrix FP framework developed in [38], as detailed below.
First, applying the matrix Lagrangian dual transform [38, Theorem 2] to the problem (11) allows us to extract the ratios from the logarithms in (9). Therefore, the problem (11) can be reformulated as
$\underset{\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}}}{\max}\ \$$f_{\rm Lag}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}}\right)$ (12a)
@@ IV-A Problem Reformulation
By applying the matrix quadratic transform [38, Theorem 1] to the problem (12), we can further decouple the ratios in the reformulated objective function (13) and reformulate the problem (12) as
$\underset{\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}}{\max}\ \$$f_{\rm Quad}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}\right)$ (14a)
@@
IV-B Update Step of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
In this step, we aim to optimize the auxiliary variables $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ with the fixed $\underline{\mathbf{W}}$, $\mathbf{T}$, and $\underline{\mathbf{R}}$. The optimal solution of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ are derived by setting the first-order derivatives of (15) to zero w.r.t. $\mathbf{\Gamma}_{k}$ and $\mathbf{\Phi}_{k}$, respectively. Let $\overline{{\mathbf{W}}}_{k}$, $\overline{\mathbf{T}}$, $\overline{\mathbf{R}}_{k}$, $\overline{\mathbf{\Gamma}}_{k}$, and $\overline{\mathbf{\Phi}}_{k}$ denote the temporal optimization results obtained by the previous iteration, and let $\overline{\mathbf{M}}_{k}$ denote the temporal interference-plus-noise matrix. The closed-form expressions for the optimal $\mathbf{\Phi}_{k}$ and $\mathbf{\Gamma}_{k}$ are given by [38]
@@ IV-B Update Step of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
IV-C Update Step of $\underline{\mathbf{W}}$
@@
$\mathbf{W}_{k}=\Big[\sum_{j=1}^{K}\sqrt{\alpha_{k}}\overline{\mathbf{H}}_{j}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\overline{\mathbf{H}}_{j}+\mu\mathbf{I}\Big]^{-1}\overline{\mathbf{H}}_{k}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right),$ (19)
where $\mu\geq 0$ is computed via bisection search to ensure that $\overline{\mathbf{W}}$ can satisfy the complementary slackness condition of the power budget constraint (11b[45].
IV-D Update Step of $\mathbf{T}$